# On $t$-structures adjacent and orthogonal to weight structures

### v1

@@ 3.3 Weight structures extended from subcategories of compact objects, and orthogonal $t$-structures
1. Then there exists a $t$-structure $t$ (left) adjacent to $w$ (on $\underline{C}$).
@@
1. According to the well-known Lemma 4.4.5 of [Nee01], the assumptions of Theorem 3.3.2 imply that $\underline{C}^{({\aleph_{0}})}$ is the retraction-closure (see §1.1) of the category ${\underline{C}^{0}}$. Note however that $w^{0}$ does not extend to $\underline{C}^{({\aleph_{0}})}$ in general (even though it extends to $\underline{C}$!); see §3.1 of [BoS18]. [^3] In any example of this sort $w^{0}$ is neither bounded above nor below; see Theorem 2.2(II.2) of ibid. Consequently, it does makes some sense not to assume ${\underline{C}^{0}}=\underline{C}^{({\aleph_{0}})}$ in our theorem.
For the same reason, the weight structures provided by our theorem don’t have to be strongly ${\aleph_{0}}$–well generated in the sense of Remark 3.3.4(1) of [Bon22a] (even if one assumes that ${\underline{C}^{0}}$ is essentially small; note that this is equivalent to the compact generation of $\underline{C}$). This observation possibly suggests that the notion of strong ${\aleph_{0}}$–well generation should be slightly generalized.
@@
Indeed, there exist lots of unbounded weight structures on (essentially) small triangulated categories; in particular, one can start from a (say, bounded) weight structure on some small non-zero ${\underline{C}^{0}}\subset\underline{C}^{({\aleph_{0}})}\subset\underline{C}$, choose a regular $\alpha>{\aleph_{0}}$, and apply Theorem 3.3.2(I.4). Next, if ${\underline{C}^{0}}=h\mathcal{C}_{0}$ is the homotopy category of a small stable $\infty$-category $\mathcal{C}_{0}$ then one can take $\underline{C}=h\operatorname{Ind}\mathcal{C}_{0}$; see Remark 1.4.4.3 of [Lur17] (cf. also the proof of Corollary 1.4.4.2 of loc. cit.). Consequently, if ${\underline{C}^{0}}=h\mathcal{C}_{0}$ is a non-zero small category endowed with a weight structure $w^{0}$ then one can extend it to an unbounded weight structure on some triangulated category $\underline{C}^{\prime}_{0}\supset{\underline{C}^{0}}$, and $\underline{C}^{\prime}_{0}\subset\underline{C}^{\prime({\aleph_{0}})}\subset\underline{C}^{\prime}$ for some compactly generated category $\underline{C}^{\prime}$.
@@
Now we recall a "geometric" setting where Proposition 3.3.4 can be applied.
@@
1. Then $\underline{C}$ is compactly generated, ${\underline{C}^{0}}=D_{perf}(X)\subset\underline{C}$ (the subcategory of perfect complexes on $X$), and $\underline{C}_{{\aleph_{0}}}=D_{coh}(\operatorname{Qcoh}(X))$; here we use the notation of Definition 2.3.4), whereas a complex $N\in\operatorname{Obj}D(\operatorname{Qcoh}(X))$ belongs to $D_{coh}(\operatorname{Qcoh}(X))$ whenever all of its cohomology sheaves $H^{i}(N)$ are coherent.
@@
The statement that $\underline{C}$ is compactly generated by $D_{perf}(X)$ is well-known; see Theorem 4.2.2(1) of [Bon24]. The calculation of $\underline{C}_{{\aleph_{0}}}$ is given by part 3 of loc. cit., and the calculation of $\underline{C}_{{\aleph_{0}}}$ is given by Proposition 4.2.9(2) of ibid. It remains to apply Proposition 2.3.5(3) to conclude the proof. ∎
@@ 3.4 $t$-structures anti-orthogonal to cosmashing weight structures
1. Then there exists a smashing $t$-structure $t^{l}$ right adjacent to $w$ and ${\underline{Ht}}^{l}$ is equivalent to the category of those additive functors ${\underline{Hw}}\to\operatorname{Ab}$ that respect products.
@@
3. Assume that $w$ is also smashing. Then for any infinite cardinal $\alpha$ the $t$-structure $t^{l}$ given by assertion 1 restricts to the subcategory $\underline{C}^{(\alpha)}$ of $\underline{C}$. Moreover, this restricted $t$-structure $t^{(\alpha)}$ is the only $t$-structure on $\underline{C}^{(\alpha)}$ that is anti-orthogonal to $w$.
@@
2. This is the easy Lemma 4.1.4 of [Nee01]. [^4] Which actually does not require any weight structure, where $\underline{C}$ is only required to be smashing.
3. The uniqueness of a $t$-structure on $\underline{C}^{(\alpha)}$ that is anti-orthogonal to $w$ is given by the dual to Corollary 2.4.4. Next, for any object $M$ of $\underline{C}^{(\alpha)}$ the functor $H^{M}$ is $\alpha$-small by the definition of $\underline{C}^{(\alpha)}$. Now, for $M^{\prime}=t^{l}_{\geq 0}M$ Proposition 2.1.2(6) says that $H^{M^{\prime}}\cong\tau_{\geq 0}H^{M}$. Hence the functor $H^{M^{\prime}}$ is $\alpha$-small as well according to Proposition 3.1.2(6), and we obtain that $M^{\prime}$ is an object of $\underline{C}^{(\alpha)}$. Thus $t^{l}$ restricts to $\underline{C}^{(\alpha)}$ indeed; see Lemma 1.1.7(1).
@@
Dualizing Theorem 3.4.1(3) one obtains that the $t$-structure $t^{r}$ provided by Theorem 3.2.3 restricts to the levels of a certain cosmallness filtration on $\underline{C}$. Yet it appears that all the levels of this filtration are zero in "reasonable cases".
@@
1. By Theorem 3.1 of [KeN13] there exists a weight structure $w$ on $\underline{C}$ such that $\underline{C}_{w\geq 0}=(\cup_{i>0}\mathcal{P}[-i]){}^{\perp}$ (i.e. $w$ is generated by $\mathcal{P}=\operatorname{Obj}\underline{A}$) and $\underline{C}_{w\leq 0}=(\cup_{i>0}\mathcal{P}[i]){}^{\perp}$. Since the category $\underline{C}$ is compactly generated by $\underline{A}$, it satisfies the dual Brown representability property (by the aforementioned Theorem 8.6.1 and Remark 6.4.5 of [Nee01]). Next, $w$ is obviously smashing and cosmashing. Applying Theorem 3.2.3(II.1), we obtain the existence of a smashing $t$-structure $t$ that is right adjacent to $w$. Since $\underline{C}_{w\leq 0}=\underline{C}_{t\leq 0}$, we obtain that $t$ is generated by $\mathcal{P}$ (as a $t$-structure) as well.
2. $t$ restricts to the subcategory $\underline{C}^{(\alpha)}$ for any infinite cardinal $\alpha$ according to Theorem 3.4.1(3). Thus it remains to prove that the classes $\underline{C}_{t\leq 0}\cap\operatorname{Obj}\underline{C}^{({\aleph_{0}})}=\underline{C}_{w\leq 0}\cap\operatorname{Obj}\underline{C}^{({\aleph_{0}})}$ and $\underline{C}_{t\geq 0}\cap\operatorname{Obj}\underline{C}^{({\aleph_{0}})}$ are the envelopes in question. The latter statement easily follow from Theorem 3.2.1(2) of [Bon22a] (see Remark 3.2.2(2) of ibid.; note also that Theorem 4.5(1,2) of [PoS16] gives this statement under the assumption that $\underline{C}$ is a "stable derivator" triangulated category). ∎
@@
    Thus we obtain a serious generalization of the existence of a (certain) $t$-structure on $\underline{C}^{({\aleph_{0}})}$ statement provided by Theorem 7.1 of [KeN13]. In particular, we do not need any differential graded algebras. Note however that our arguments do not give the description of $t^{{\aleph_{0}}}$ in terms of "generators" that would be similar to that in loc. cit.
@@
    Since $t^{\prime}$ is bounded and ${\underline{Ht}}^{\prime}$ is a length category, the category $\underline{C}^{({\aleph_{0}})}$ is densely generated by $\operatorname{Obj}\underline{A}$; hence $\underline{A}$ is easily seen to generate $\underline{C}$ (in the sense of Definition 3.2.1(3)). Thus one can apply Corollary 3.4.3 to this setting. Moreover, it is easily seen that our assumptions on $t^{\prime}$ (combined with part 2 of our corollary) imply that the corresponding $t$-structure $t^{{\aleph_{0}}}$ coincides with $t^{\prime}$.
  4. 4.
    It would be interesting to find which assumptions on a general $t$-structure $t^{\prime}$ on $\underline{C}^{({\aleph_{0}})}$ ensure that $t^{\prime}$ extends to a $t$-structure on $\underline{C}$ and there exists a weight structure $w$ that is right adjacent to $t$.
4 On functors and $t$-structures related to (essentially) bounded weight structures
@@
To formulate our statements in a more general case we recall (in §4.1) the notions of essentially bounded (above, below, or both) objects and weight structures; we motivate this "almost generalization" in Remarks 4.1.5(2,3), 4.2.3(2), and 4.2.7 below. We also study weight structures of these types as well as virtual $t$-truncations with respect to them and their relation with certain "essential weight-boundedness" for functors.
In §4.2 we apply the aforementioned results to the construction of certain adjacent and orthogonal $t$-structures. We apply our general statements to various derived categories of (quasi)coherent sheaves. In particular, we are able to generalize some central results of [Bon24].
4.1 On essentially bounded weight structures and the relation to cohomological functors
@@
5. A cohomological functor $H$ from $\underline{C}$ will be said to be locally bounded below (resp. above) [^5] This terminology was inspired by [Nee18, Remark 0.2]. if for any $M\in\operatorname{Obj}\underline{C}$ we have $H^{i}(M)=0$ for all $i\ll 0$ (resp. $i\gg 0$).
@@ 1. Assume first that $M$ is essentially $w$-bounded below and there exist distinguished triangles
We can assume that $n_{1}\geq n_{2}$. Then both $T_{1}$ and $T_{2}$ are $m$-weight decompositions of $\tilde{M}$ for any $n<n_{2}$. Applying the uniqueness statement in Proposition 1.2.4(4) one can easily obtain $T_{1}\cong T_{2}$. In particular, one can note that both $T_{1}$ and $T_{2}$ give weight decompositions of $\tilde{M}$ that avoid weight $n_{2}-1$; see Theorem 2.2.1(9) of [Bon22b]. Now we applying loc. cit. to obtain $T_{1}\cong T_{2}$.
The case of a $w$-bounded above $M$ is just the categorical dual of assertion 1; see Proposition 1.2.4(1).
2. One may say that this is a "one-sided unbounded version of" [BoV23, Theorem 2.5(I)].
Firstly, conditions 1 and 2 are equivalent by Corollary 3.1.4 (resp. Theorem 3.1.3) of [Bon22b]. Moreover, these statements also yield that condition 2 follows from 3; recall here that pure functors (cf. condition 7 of Corollary 3.1.4 and condition 8 of Theorem 3.1.3 in ibid.) are the functors of weight range $[0,0]$. [^6] Actually, in loc. cit. mentions homological pure functors $\underline{C}\to\underline{A}$; yet one can just reverse the arrows in the target category. The converse implication is valid as well; see conditions 2 and Theorem 2.5(I) of [BoV23].
@@
1. Assume that $M[-n]$ is essentially $w$-positive (resp. $w$-negative, for some $n\in{\mathbb{Z}}$), and $m<n-1$ (resp. $m>n+1$) and $H$ is a cohomological functor from $\underline{C}$. Then $\tau_{\leq m}(H)(\tilde{M})\cong H(RD(\tilde{M}))$ (resp. $\tau_{\geq m}(H)(\tilde{M})\cong H(LD(\tilde{M}))$). Consequently, $\tau_{\leq m}(H)(\tilde{M})$ (resp. $\tau_{\geq m}(H)(\tilde{M})$) is a retract of $H(RD(\tilde{M}))$ (resp. of $H(L(RD(\tilde{M})))$).
@@
If $w$ is weight-Karoubian, that is, if ${\underline{Hw}}$ is Karoubian then one modify the definition of essentially $w$-bounded above and below objects by setting $\tilde{M}=M$ in the triangles (4.1.1); see Theorems 2.3.4 and 3.1.3 and Corollary 3.1.4 of [Bon22b]. Respectively, one can compute $\tau_{\leq m}(H)(M)$ for any $M\in\operatorname{Obj}\underline{C}$ using the corresponding canonical decomposition similarly to Proposition 4.1.4(1). 2. Obviously, any semi-orthogonal decomposition couple is an essentially bounded weight structure; see condition 4 of Proposition 1.2.6. Furthermore, one may say that general essentially bounded above and below weight structures are "mixes" of bounded above and below ones and semi-orthogonal decompositions. Indeed, if $w$ is weight-Karoubian and essentially bounded above or below then the statements mentioned in part 1 of this remark yield the existence of a certain semi-orthogonal decomposition. 3. Respectively, the main results of this section can be used to generalize some statement on semi-orthogonal decompositions from [Bon24]. We will say more one this in Remark 4.2.7(2) below.
4.2 More on the construction of $t$-structures
Now we pass to the existence of $t$-structures orthogonal to essentially bounded weight structures. We start with a nice "bounded $R$-linear analogue" of Theorem 3.2.3. Once again, we assume that $R$ is a commutative Noetherian ring).

### v2

@@ 3.3 Weight structures extended from subcategories of compact objects, and orthogonal $t$-structures
1. Then there exists a $t$-structure $t$ (left) adjacent to $w$.
@@
Indeed, there exist lots of unbounded weight structures on (essentially) small triangulated categories; in particular, one can start from a (say, bounded) weight structure on some small non-zero ${\underline{C}^{0}}\subset\underline{C}^{({\aleph_{0}})}\subset\underline{C}$, choose a regular $\alpha>{\aleph_{0}}$, and apply Theorem 3.3.2(I.4) on the corresponding category $\underline{C}_{\alpha}^{0}\subset\underline{C}$ (that is the smallest subcategory of $\underline{C}$ that is closed with respect to coproducts of less than $\alpha$ objects and contains ${\underline{C}^{0}}$). Next, if ${\underline{C}^{0}}=h\mathcal{C}_{0}$ is the homotopy category of a small stable $\infty$-category $\mathcal{C}_{0}$ then it compactly generates the category $\underline{C}=h\operatorname{Ind}\mathcal{C}_{0}$; see Remark 1.4.4.3 of [Lur17] (cf. also the proof of Corollary 1.4.4.2 of loc. cit.).
Consequently, if ${\underline{C}^{0}}=h\mathcal{C}_{0}$ is a non-zero small category endowed with a weight structure $w^{0}$ then one can extend it to an unbounded weight structure on some triangulated category $\underline{C}^{\prime}_{0}\supset{\underline{C}^{0}}$, and $\underline{C}^{\prime}_{0}\subset\underline{C}^{\prime({\aleph_{0}})}\subset\underline{C}^{\prime}$ for some compactly generated category $\underline{C}^{\prime}$.
@@
Now we recall a ’geometric’ setting where Proposition 3.3.4 can be applied.
@@
1. Then $\underline{C}$ is compactly generated, ${\underline{C}^{0}}=D_{perf}(X)\subset\underline{C}$ (the subcategory of perfect complexes on $X$), and $\underline{C}_{{\aleph_{0}}}=D_{coh}(\operatorname{Qcoh}(X))$; here we use the notation of Definition 2.3.4, whereas a complex $N\in\operatorname{Obj}D(\operatorname{Qcoh}(X))$ belongs to $D_{coh}(\operatorname{Qcoh}(X))$ whenever all of its cohomology sheaves $H^{i}(N)$ are coherent.
@@
The statement that $\underline{C}$ is compactly generated by $D_{perf}(X)$ is well-known; see Theorem 4.7(1) of [Bon24b]. The calculation of $\underline{C}_{{\aleph_{0}}}$ is given by part 3 of loc. cit., and the calculation of $\underline{C}_{\underline{A}^{T}}$ is given by Proposition 4.14(2) of ibid. It remains to apply Proposition 2.3.5(3) to conclude the proof. ∎
@@ 3.4 $t$-structures anti-orthogonal to cosmashing weight structures
1. Then there exists a smashing $t$-structure $t^{r}$ right adjacent to $w$ and ${\underline{Ht}}^{r}$ is equivalent to the category of those additive functors ${\underline{Hw}}\to\operatorname{Ab}$ that respect products.
@@
3. Assume that $w$ is also smashing. Then for any infinite cardinal $\alpha$ the $t$-structure $t^{r}$ given by assertion 1 restricts to the subcategory $\underline{C}^{(\alpha)}$ of $\underline{C}$. Moreover, this restricted $t$-structure $t^{(\alpha)}$ is the only $t$-structure on $\underline{C}^{(\alpha)}$ that is anti-orthogonal to $w$.
@@
2. This is the easy Lemma 4.1.4 of [Nee01] (if one ignores $w$).
3. The uniqueness of a $t$-structure on $\underline{C}^{(\alpha)}$ that is anti-orthogonal to $w$ is given by the dual to Corollary 2.4.4. Next, for any object $M$ of $\underline{C}^{(\alpha)}$ the functor $H^{M}$ is $\alpha$-small by the definition of $\underline{C}^{(\alpha)}$. Now, for $M^{\prime}=t^{r}_{\geq 0}M$ Proposition 2.1.2(6) says that $H^{M^{\prime}}\cong\tau_{\geq 0}H^{M}$. Hence the functor $H^{M^{\prime}}$ is $\alpha$-small as well according to Proposition 3.1.2(6), and we obtain that $M^{\prime}$ is an object of $\underline{C}^{(\alpha)}$. Thus $t^{r}$ restricts to $\underline{C}^{(\alpha)}$ indeed; see Lemma 1.1.5(1).
@@
Dualizing Theorem 3.4.1(3) one obtains that the $t$-structure $t^{r}$ provided by Theorem 3.2.3 restricts to the levels of a certain cosmallness filtration on $\underline{C}$. Yet it appears that all the levels of this filtration are zero in ’reasonable cases’.
@@
1. By Theorem 3.1 of [KeN13], there exists a weight structure $w$ on $\underline{C}$ such that $\underline{C}_{w\geq 0}=(\cup_{i>0}\mathcal{P}[-i]){}^{\perp}$ (i.e. $w$ is generated by $\mathcal{P}=\operatorname{Obj}\underline{A}$) and $\underline{C}_{w\leq 0}=(\cup_{i>0}\mathcal{P}[i]){}^{\perp}$. Since the category $\underline{C}$ is compactly generated by $\underline{A}$, it satisfies the dual Brown representability property (by the aforementioned Theorem 8.6.1 and Remark 6.4.5 of [Nee01]). Next, $w$ is obviously smashing and cosmashing. Applying Theorem 3.2.3(II.1), we obtain the existence of a smashing $t$-structure $t$ that is right adjacent to $w$. Since $\underline{C}_{w\leq 0}=\underline{C}_{t\leq 0}$, we obtain that $t$ is generated by $\mathcal{P}$ (as a $t$-structure) as well.
2. $t$ restricts to the subcategory $\underline{C}^{(\alpha)}$ for any infinite cardinal $\alpha$ according to Theorem 3.4.1(3). Thus it remains to prove that the classes $\underline{C}_{t\leq 0}\cap\operatorname{Obj}\underline{C}^{({\aleph_{0}})}=\underline{C}_{w\leq 0}\cap\operatorname{Obj}\underline{C}^{({\aleph_{0}})}$ and $\underline{C}_{t\geq 0}\cap\operatorname{Obj}\underline{C}^{({\aleph_{0}})}$ are the envelopes in question. The latter statement easily follow from Theorem 3.2.1(2) of [Bon22] (see Remark 3.2.2(2) of ibid.). [^4] Note also that Theorem 4.5(1,2) of [PoS16] gives this statement under the assumption that $\underline{C}$ is a ’stable derivator’ triangulated category.
@@
    Thus we obtain a serious generalization of the existence of a (certain) $t$-structure on $\underline{C}^{({\aleph_{0}})}$ statement provided by Theorem 7.1 of [KeN13]. In particular, we do not need any differential graded algebras (and can construct stable $\infty$-exampes for Corollary 3.4.3; cf. Remark 3.3.3(2)). Note however that our arguments do not give the description of $t^{{\aleph_{0}}}$ in terms of ’generators’ that would be similar to that in loc. cit.
@@
    Since $t^{\prime}$ is bounded and ${\underline{Ht}}^{\prime}$ is a length category, the category $\underline{C}^{({\aleph_{0}})}$ is densely generated (see §1.1) by $\operatorname{Obj}\underline{A}$; hence $\underline{A}$ is easily seen to generate $\underline{C}$ (in the sense of Definition 3.2.1(3)). Thus one can apply Corollary 3.4.3 to this setting. Moreover, it is easily seen that our assumptions on $t^{\prime}$ (combined with part 2 of our corollary) imply that the corresponding $t$-structure $t^{{\aleph_{0}}}$ coincides with $t^{\prime}$.
4 On functors and $t$-structures related to essentially bounded weight structures
@@
To formulate our statements in a more general case we recall (in §4.1) the notions of essentially bounded (above, below, or both) objects and weight structures; we motivate these definitions in Remarks 4.1.5(2,3), 4.2.3(2), and 4.2.7 below. We also study weight structures of these types as well as virtual $t$-truncations with respect to them and their relation with certain essential weight-boundedness for functors.
In §4.2 we apply the aforementioned results to the construction of certain adjacent and orthogonal $t$-structures. We apply our general statements to various derived categories of (quasi)coherent sheaves. In particular, we are able to generalize some central results of [Bon24b].
4.1 Essentially bounded weight structures and their relation to cohomological functors
@@
5. A cohomological functor $H$ from $\underline{C}$ will be said to be locally bounded below (resp. above) [^5] This terminology was inspired by Remark 0.2 of [Nee18]. if for any $M\in\operatorname{Obj}\underline{C}$ we have $H^{i}(M)=0$ for all $i\ll 0$ (resp. $i\gg 0$).
@@ 1. Assume first that $M$ is essentially $w$-bounded below and there exist distinguished triangles
We can assume that $n_{1}\geq n_{2}$. Then both $T_{1}$ and $T_{2}$ are $m$-weight decompositions of $\tilde{M}$ for any $n<n_{2}$. Applying the uniqueness statement in Proposition 1.2.4(4) one can easily obtain $T_{1}\cong T_{2}$. In particular, one can note that both $T_{1}$ and $T_{2}$ give weight decompositions of $\tilde{M}$ that avoid weight $n_{2}-1$; see Theorem 2.2.1(9) of [Bon24a]. Now we apply loc. cit. to obtain $T_{1}\cong T_{2}$.
The case of an essentially $w$-bounded above $M$ is just the categorical dual of the essentially bounded below one; see Proposition 1.2.4(1).
2. One may say that this is a one-sided unbounded analogue of [BoV23, Theorem 2.5(I)].
Firstly, conditions 1 and 2 are equivalent by Corollary 3.1.4 (resp. Theorem 3.1.3) of [Bon24a]. Moreover, these statements also yield that condition 2 follows from 3; recall here that pure functors (cf. condition 7 of Corollary 3.1.4 and condition 8 of Theorem 3.1.3 in ibid.) are the functors of weight range $[0,0]$. [^6] Actually, in loc. cit. mentions homological pure functors $\underline{C}\to\underline{A}$; yet one can just reverse the arrows in the target category. The converse implication is valid as well; see conditions 2 and Theorem 2.5(I) of [BoV23].
@@
1. Assume that $M[-n]$ is essentially $w$-positive (resp. $w$-negative, for some $n\in{\mathbb{Z}}$), $m<n-1$ (resp. $m>n+1$), and $H$ is a cohomological functor from $\underline{C}$. Then $\tau_{\leq m}(H)(\tilde{M})\cong H(RD(\tilde{M}))$ (resp. $\tau_{\geq m}(H)(\tilde{M})\cong H(LD(\tilde{M}))$). Consequently, $\tau_{\leq m}(H)(M)$ (resp. $\tau_{\geq m}(H)(M)$ ) is a retract of $H(RD(\tilde{M}))$ (resp. of $H(LD(\tilde{M}))$).
@@
If $w$ is weight-Karoubian, that is, if ${\underline{Hw}}$ is Karoubian then one modify the definition of essentially $w$-bounded above and below objects by setting $\tilde{M}=M$ in the triangles (4.1.1); see Theorems 2.3.4 and 3.1.3 and Corollary 3.1.4 of [Bon24a]. Respectively, one can compute $\tau_{\leq m}(H)(M)$ for any $M\in\operatorname{Obj}\underline{C}$ using the corresponding canonical decomposition similarly to Proposition 4.1.4(1).
2. Obviously, any semi-orthogonal decomposition couple is an essentially bounded weight structure (if we permute the components); see condition 4 of Proposition 1.2.6.
Furthermore, one may say that general essentially bounded above and below weight structures are ’mixes’ of bounded above and below ones and semi-orthogonal decompositions. Indeed, if $w$ is weight-Karoubian and essentially bounded above or below then the statements mentioned in part 1 of this remark yield the existence of a certain semi-orthogonal decomposition.
3. Respectively, the main results of this section can be used to generalize some statement on semi-orthogonal decompositions from [Bon24b]. We will say more one this in Remark 4.2.7(2) below.
4.2 On $t$-structures corresponding to essentially bounded weight structures
Now we pass to the existence of $t$-structures orthogonal to essentially bounded weight structures. We start with a nice ’bounded $R$-linear analogue’ of Theorem 3.2.3. Once again, we assume that $R$ is a commutative Noetherian ring.